What is the boiling point elevation of a solution that contains 40.5 grams of glycerol in 100 g of water?



1
Expert's answer
2021-01-25T04:23:28-0500
"\Delta T=E*Cm"

"\Delta T" - boiling point elevation; K

E - Ebulioscopic constant for water; E=0.52 K*kg/mol

Cm - mol/kg.


"Cm=m*100\/(M*m(H2O))"

"Cm=40.5*1000\/92\/100=4.4 mol\/kg"

"\Delta T=0.52*4.4=2.288 K"